OpenStudy (anonymous):

Simplify a * 3b^9 * 6a

OpenStudy (shamil98):

what is 6a * a ?

OpenStudy (shamil98):

do that first. and i'll move on to the next step.

OpenStudy (anonymous):

multiplication

OpenStudy (mathstudent55):

With the commutative and associative properties, you can change the order of the operations and the grouping of the factors. \(a \times 3b^9 \times 6a \) \(= 3 \times 6 \times a \times a \times b^9\)

OpenStudy (shamil98):

\[\huge 6a * a * 3b^9 \] multiply \[\huge (6*3)(a*a)(b^9)\] can you do it like this?

OpenStudy (anonymous):

yeah